OpenStudy (anonymous):

Solve. m-5 = 1/35 (m to the power of -5)

OpenStudy (hockeychick23):

m-5+5= 1/35 + 5 m= 5 1/35

OpenStudy (anonymous):

its m to the power of -5

OpenStudy (anonymous):

\[\frac{ 1 }{ m^5 }=\frac{ 1 }{ 32 }\]

OpenStudy (anonymous):

\[m^5=32\]

OpenStudy (anonymous):

\[m=\sqrt[5]{32}\]

OpenStudy (anonymous):

\[m=32^{\frac{ 1 }{ 5 }}\]

OpenStudy (anonymous):

\[M=(2^5)^{\frac{ 1 }{ 5 }}\]\[m=?\]

OpenStudy (anonymous):

solve it @Bilaal.K

OpenStudy (anonymous):

i dont really know what to do next

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

@Somy @Preetha @dan815 @shaik0124 @ayyookyndall @cwrw238 @caitlinnr14 @lorrsow40 @hockeychick23 @Danielle05 @sweeta @Kainui @kohai @iGreen @the_cool_moose

OpenStudy (hockeychick23):

your answer is 2

OpenStudy (anonymous):

The answer is 2?

OpenStudy (hockeychick23):

yes

OpenStudy (anonymous):

ok can you please help me with some more?